OpenStudy (anonymous):

There are a total of 15 apartments in two buildings. The difference of two times the number of apartments in the first building and three times the number of apartments in the second building is 5. a. Write a system of equations to model the relationship between the number of apartments in the first building f and the number of apartments in the second building s. b. How many apartments are in each building? (if it helps this deals with elimination)

OpenStudy (anonymous):

15 total 15= f+s 2f-3s =5 solve: f+s = 15 2f-3s =5 you can solve this various ways... you can multiply the top eqn by 3 and add both eqns, so you have 3f + 3s = 45 2f-3s =5 ----------- 5f = 50 f=10, plug f=10 into one of the eqns and get s

OpenStudy (anonymous):

thank you person ^^ that's all i really needed i just get confuzzled with word problems

OpenStudy (anonymous):

okay. Great!